An empty freight car A of mA = 2.50 x 104kg coasts along a horizontal track at 2.70 m/s until it couples to a stationary car B of mB = 4.75 x 104kg. There is negligible friction with the track, and the brakes are off. What is the speed of the two cars as they move along the track after the interaction?



1
Expert's answer
2020-12-28T08:41:56-0500
"m_AV=(m_A+m_B)V'\\(2.5\cdot10^4)2.7=(2.5\cdot10^4+4.75\cdot10^4)V'\\V'=0.931\frac{m}{s}"